OpenStudy (anonymous):

The sum of two consecutive intergers is 153. Which equation can be used to find the first interger n? 2n+1=153 2n=153 n+1=153 2n+2=153

OpenStudy (radar):

If we let the first integer be equal to n, how would we represent the next consecutive integer?

OpenStudy (anonymous):

Idk

OpenStudy (radar):

What does consecutive mean? which pairs would you consider being consecutive? 23, 24 30, 32 3,4 71, 72 ??

OpenStudy (anonymous):

Idk could you explain this to me like I am an 5 year old please.

OpenStudy (radar):

two consecutive integers means the numbers are next to each other on the number line. In the above, 23 and 24 are consecutive so are 3 and 4, do you see another pair above that was consecutive?

OpenStudy (anonymous):

Maybe you could use the problem as your example. Don't give me the answer just explain it to me.

OpenStudy (radar):

O.K. Do you know algebra?

OpenStudy (radar):

Here is the problem: The sum of two consecutive intergers is 153. Which equation can be used to find the first interger n? 2n+1=153 2n=153 n+1=153 2n+2=153 Using the problem as an example: Let n = the first (are smaller) number. Then we would state that n+1 = the next consecutive number. These two numbers sum up to 153 so we write. n + n + 1 = 153. consolidating similar terms we get 2n + 1 = 153.

OpenStudy (radar):

*are should be or